### 09 Jan 2024 v1

# Reconnaissance with JWST of the J-region Asymptotic Giant Branch in Distance Ladder Galaxies: From Irregular Luminosity Functions to Approximation of the Hubble Constant

Abstract
We study stars in the J-regions of the asymptotic giant branch (JAGB) of near-infrared color magnitude diagrams in the maser host NGC 4258 and 4 hosts of 6 Type Ia supernovae (SN Ia): NGC 1448, NGC 1559, NGC 5584, and NGC 5643. These clumps of stars are readily apparent near and =22-25 mag with James Webb Space Telescope NIRCam photometry. Various methods have been proposed to assign an apparent reference magnitude for this recently proposed standard candle, including the mode, median, sigma-clipped mean or a modeled parameter of its luminosity function. We test the consistency of these by measuring intra-host variations, finding differences of up to 0.2 mag that significantly exceed statistical uncertainties. Brightness differences appear intrinsic, and are further amplified by the non-uniform shape of the JAGB LF, an issue already apparent in comparing the LMC and SMC. We follow a “many methods” approach to consistently measure JAGB magnitudes and distances to the SN Ia host sample calibrated by NGC 4258. We find broad agreement with the distances measured from Cepheids, tip of the red giant branch (TRGB) and Miras. However, the SN host mean distance estimated via the JAGB method and necessary to estimate , differs by 0.18 mag amongst the above definitions, a result of the different levels of asymmetry of the JAGB luminosity functions. The methods yield a full range of km s Mpc, i.e., a fiducial result of (stat) 2.3 (sys) ( 3.2 if combined in quadrature) km s Mpc, independent of Cepheids, TRGB, or Miras, with systematic errors limited by the differences in methods. Future work may seek to further standardize and refine this promising tool, making it more competitive with established distance indicators.
1 Introduction
The need for primary distance indicators that can reach tens of Mpc motivates efforts to identify alternative luminous standard candles, such as the recently proposed stars in the J-region Asymptotic Giant Branch (JAGB). The JAGB can be identified as an enhanced density of stars or a “clump” in the near-infrared color magnitude diagram (CMD; often in vs. or a similar filter set) redward of the red giant branch (RGB) and brighter than the tip of the RGB . This region is labelled ‘J’ in Fig. 1, corresponding to the labels defined by Weinberg & Nikolaev (2001), and is expected from stellar theory to be populated by thermally-pulsing carbon-rich (photospheric carbon-to-oxygen ratio ), dust-producing asymptotic giant branch stars with estimated amplitudes of 0.7 mag (Madore & Freedman, 2020) and J-band luminosities of 6 mag. Various measures of the apparent magnitude of the JAGB region, such as the mode, mean, median or a fitted value of its luminosity function, have been suggested and used in the literature as a potential standard candle that is luminous enough to measure extragalactic distances at megaparsec scales (Battinelli & Demers, 2005; Madore & Freedman, 2020; Freedman & Madore, 2020; Ripoche et al., 2020; Lee et al., 2021b; Zgirski et al., 2021; Parada et al., 2021; Madore et al., 2022; Parada et al., 2023; Lee, 2023).
The theoretical foundation for expecting the mean of the JAGB to be a useful standard candle is rooted in the prediction that only a narrow range of stellar masses allows for a large quantity of carbon to be transported to the star’s surface, which results in the formation of carbonaceous circumstellar dust. These stars are expected to be of intermediate age and metallicity, with ages ranging from 300 Myr to 1 Gyr and masses ranging from 2 to 5 (Marigo et al., 2008; Karakas, 2014; Madore & Freedman, 2020), although these ranges may change with metallicity of the stars (Karakas, 2014). When a star’s mass falls within the prescribed range, the third dredge-up events during post-main sequence mass transport lead to the movement of carbon to the surface, resulting in the formation of a carbon AGB star (Groenewegen & Marigo, 2004). Stars with too little mass () do not have dredge up events that are efficient enough to bring enough carbon to the surface of the star to form a carbon star. Stars with too much mass will experience hot bottom burning, where carbon burns at the bottom of the convection zone before it can reach the surface. Observationally, the C/O1 feature of carbon stars is characteristically shown by carbon spectral bands such as CN, C, and CO, and lack of oxides (see for instance Gonneau et al., 2017, and references therein).
Previous studies have indicated that the mass range required for an oxygen-rich asymptotic giant branch (AGB) star to transition into a carbon-rich AGB star can vary with the metallicity of the stellar population (Karakas, 2014; Ripoche et al., 2020; Parada et al., 2021, 2023). Lee (2023) showed that the JAGB does not show any dependency of its luminosity with metallicity over a very small explored range in [M/H] of 0.08 dex (). However, due to the small range in metallicity studied, we caution that their data is simultaneously consistent with no trend and with what could be a very large trend of 1 mag/dex. To show this, we show in Fig. 2 the data points from the reddening corrected JAGBs in Fig. 8 of Lee (2023); the nominal best-fit trend (red line) has a slope of mag/dex.111We also show the Spearman’s coefficient in the bottom left hand corner, which matches the value found in Lee (2023). The limited leverage due to the small range of metallicity results in a constraint that is consistent to within 1- with both no metallicity dependence and what would be a very large metallicity dependence of 1 mag/dex (cyan line in Fig. 2). For reference, we also plot a line with slope of 0.2 mag/dex, corresponding to approximately the most widely found Cepheid metallicity dependence (Breuval et al., 2022; Trentin et al., 2023, and references therein). All three of the lines shown in Fig. 2 are consistent with the data presented in Lee (2023), indicating no strong constraint. For context, if the metallicity dependence was as large as 1 mag/dex, the effort to calibrate to 5% precision without any metallicity correction would restrict the measurement to a field about 0.3 arcminutes wide at the distances of SN Ia host galaxies, based on typical host metallicity gradients of 0.05 dex/kpc (Hoffmann et al., 2016).
Thus it would be risky to assume the JAGB is not impacted by differences in host population metallicity (or age) at a level needed to support a few-percent-level determination of the Hubble constant. However, few studies have yet quantified the potential size of such intrinsic variations in the JAGB candle at the target scale (Madore & Freedman, 2020; Freedman & Madore, 2020; Lee et al., 2021b, a; Zgirski et al., 2021; Lee et al., 2022, 2021a; Lee, 2023). Karakas (2014) used theoretical stellar evolutionary models of AGB stars to demonstrate that higher metallicities in stellar populations lead to less efficient dredge-up events.
Another issue which challenges the use of JAGB as a new standard candle derives from the change in shape (or population mix) indicated in its luminosity function between different hosts. Parada et al. (2021) show significantly different levels of asymmetry in the JAGB LFs in different galaxies: the LMC and NGC 6822 exhibit highly skewed luminosity functions, in contrast to the symmetric luminosity functions in the SMC and IC 1613; these differences are not due to foreground or background contamination. In the Local Group, the asymmetric LF in the LMC results produces a difference between its mode and mean of 0.15 mag, while this difference for the SMC is only mag. The magnitude estimates based on the median or a tightly-framed mean are in good agreement with their relative geometric distances; the distances based on the mode are not (Parada et al., 2021). We will revisit this point in the Discussion. Tight-framing of the mean (i.e., the selection of a small range to average), as employed by Madore & Freedman (2020) and Freedman & Madore (2020), can be used to avoid population contamination, and may offer good agreement with other distance indicators, but the lack of a defined prescription for these tight frames makes it difficult to evaluate and replicate this procedure. The non-trivial width of the J-region LF ( 0.3 mag for a single epoch and 0.2 mag using time-averaged magnitudes) also suggests some diversity of the stellar populations in this region. Despite these issues, the JAGB LF still offers one of the few possible standard candles luminous enough to rival the distance reach of Cepheids, TRGB, and Miras (Lee et al., 2023), motivating the present empirical study.
We find the JAGB clumps are readily apparent in the CMDs measured from JWST Cycle 1 program (GO-1685, P.I.: A. Riess, Riess et al., 2021), which observed four SN Ia hosts within 25 Mpc and NGC 4258 (obtained to measure Cepheids, Miras, and the TRGB). This further motivated us to investigate the feasibility of using the JAGB as a standard candle following the methods employed in the literature. We first investigate variations in the measured JAGB within hosts. We then measure distances and following a distance ladder and “many methods” approach. In Sections 2.1 to 2.3 we describe the data processing used for this study. We introduce how we measure the JAGB in Section 2.4. In Section 3 and 3.3, we investigate the JAGB through null tests, by varying the region of the measurement within a host. We characterize the variations in the JAGB measurements and then use them to determine distances to the SN Ia hosts in Section 4. We discuss our results in Section 5.
2 Measurement
2.1 Data Selection
We retrieved F090W (0.9 m), F150W (1.5 m), and F277W (2.8 m) images of NGC 4258, NGC 1448, NGC 1559, NGC 5584, and NGC 5643 from Cycle 1 program GO-1685 (P.I.: A. Riess, Riess et al., 2021) from the Mikulski Archive for Space Telescopes222https://archive.stsci.edu/. Recently, STScI implemented a major update to their image pipeline with new flat fields and zeropoints in the jwst_1125.pmap and jwst_1126.pmap context files. After the release of jwst_1126.pmap, there have been no major updates that would significantly affect NIRCam photometry. For this study, we use images processed with context files released later than jwst_1126.pmap. The versions of these context files can be found in Table 1.
The primary purpose of these observations was to observe a large sample of Cepheids in the disks and TRGB in the halos of the anchor galaxy, NGC 4258, as well as multiple SN Ia host galaxies to construct a distance ladder to measure . Consequently, these observations contain two visits, visit 1 and visit 2. Because the JAGB is measured in the outer disk (regions younger than old halos, hence with a population of AGB stars), this program necessarily samples regions between the sites of Cepheids and the TRGB.
We provide a summary of the observations analyzed in this study in Table 1 and show their footprints in Fig. 3. For all JAGB measurements presented below, we correct for foreground extinction by retrieving values using the dustmaps package (Green, 2018) in Python and galaxy coordinates from the NASA/IPAC Extragalactic Distance Database (NED)333https://ned.ipac.caltech.edu/. The NASA/IPAC Extragalactic Database (NED) is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology. and applying the extinction coefficients listed in Table 3 assuming =3.1 and = 0.6021, and 0.2527 for F150W, and F277W, respectively. We measure all JAGBs in the main body of this paper in F150W.
| Galaxy | Observation | Observation Date | Filters | Exposure Time [s] | Context File Number |
| NGC 1448 | 13 | 2023-08-02 | F090W/F277W | 418.7 x 4 | 1146 |
| NGC 1448 | 13 | 2023-08-02 | F150W/F277W | 526.1 x 4 | ” |
| NGC 1448 | 14 | 2023-08-18 | F090W/F277W | 418.7 x 4 | ” |
| NGC 1448 | 14 | 2023-08-18 | F150W/F277W | 526.1 x 4 | ” |
| NGC 1559 | 1 | 2023-06-30 | F090W/F277W | 418.7 x 4 | ” |
| NGC 1559 | 1 | 2023-06-30 | F150W/F277W | 526.1 x 4 | ” |
| NGC 1559 | 2 | 2023-07-15 | F090W/F277W | 418.7 x 4 | ” |
| NGC 1559 | 2 | 2023-07-15 | F150W/F277W | 526.1 x 4 | ” |
| NGC 4258 | 5 | 2023-05-02 | F090W/F277W | 257.7 x 4 | 1147 |
| NGC 4258 | 5 | 2023-05-02 | F150W/F277W | 365.1 x 4 | ” |
| NGC 4258 | 6 | 2023-05-17 | F090W/F277W | 257.7 x 4 | ” |
| NGC 4258 | 6 | 2023-05-17 | F150W/F277W | 365.1 x 4 | ” |
| NGC 5584 | 9 | 2023-01-30 | F090W/F277W | 418.7 x 4 | 1132 |
| NGC 5584 | 9 | 2023-01-30 | F150W/F277W | 526.1 x 4 | ” |
| NGC 5584 | 10 | 2023-02-21 | F090W/F277W | 418.7 x 4 | ” |
| NGC 5584 | 10 | 2023-02-21 | F150W/F277W | 526.1 x 4 | ” |
| NGC 5643 | 11 | 2023-07-07 | F090W/F277W | 311.4 x 4 | 1146 |
| NGC 5643 | 11 | 2023-07-07 | F150W/F277W | 418.7 x 4 | ” |
| NGC 5643 | 12 | 2023-07-22 | F090W/F277W | 311.4 x 4 | ” |
| NGC 5643 | 12 | 2023-07-22 | F150W/F277W | 418.7 x 4 | ” |
| Galaxy | RA | Dec | E(B-V) | ||
| NGC 4258 | 184.73958 | 47.30397 | 0.016 | 0.010 | 0.004 |
| NGC 1448 | 56.13300 | 44.64483 | 0.014 | 0.009 | 0.004 |
| NGC 1559 | 64.39904 | 62.78367 | 0.030 | 0.018 | 0.008 |
| NGC 5584 | 215.59904 | 0.38767 | 0.039 | 0.024 | 0.010 |
| NGC 5643 | 218.16975 | 44.17442 | 0.169 | 0.102 | 0.043 |
2.2 DOLPHOT Photometry
We measure PSF photometry using the NIRCam version of DOLPHOT (Dolphin, 2000, 2016) in the Vega-Vega system. For this analysis, we used the beta version of the NIRCam module of DOLPHOT released on April 6, 2023. The most recent update as of writing is from December 2, 2023; this includes two changes since the April 6, 2023 version that we do not expect to significantly affect our results. The first change was the inclusion of the ‘Sirius-Vega’ zero-point. In this paper, we choose to remain with the original ‘Vega-Vega’ zero-point. The second change incorporates the “etctime” option, which allows the user to adjust the exposure times provided in the image headers. The choice of exposure times do not affect the photometry, although it may affect the photometric errors estimated with Poisson statistics. The photometric errors for our observations are in general too small to have a significant impact on the JAGB luminosity function, which has a typical width of 0.2 mag or more. See also Anand et al. (in prep.) for a discussion of these features. In addition, any future updates will likely minimally affect the results presented here because our measurements are all relative (i.e., between SN Ia hosts and NGC 4258 or within a host).
We use stage 3 F150W *i2d.fits frames as reference frames, which are created from combining multiple visit images, and perform photometry on stage 2 *cal.fits frames, which are the single frames corresponding to each dither. We use the DOLPHOT parameters recommended in the NIRCam DOLPHOT manual444http://americano.dolphinsim.com/dolphot/dolphotNIRCam.pdf and apply the same quality cuts to the F090W and F150W sample as used by Riess et al. (2023), which were modified from Warfield et al. (2023): 1) Crowding 0.5, 2) Sharpness 0.01, 3) Object Type 2, 4) S/N 3, 5) Error Flag 2. As described in Riess et al. (2023), there is currently an issue in the NIRCam DOLPHOT source detections when long wavelength images (F277W) are combined with with short wavelength images (F090W, F150W) when performing photometry. To circumvent this issue, we perform photometry using the “warmstart” option in DOLPHOT and perform a first round of photometry on the short wavelength images (F090W and F150W), then use the resulting source list to find stars in a second round of photometry that includes all three filters. For our final measurements, we use the F150W photometry from the first run and the F277W from the second warmstart run (see Riess et al. 2023 for a detailed explanation of this setup).
2.3 Crowding Corrections
Because of the moderate stellar density present in the outer disks where we measured the JAGB, we estimated the impact of crowding on photometry by measuring “crowding corrections”. We derive and apply crowding corrections for all stars located inside the outer disk fields listed in Table 5 using DOLPHOT artificial star measurements. To determine these crowding corrections, we divide each image in these disk fields into a 10x10 grid. Next, we place 200 artificial stars in each grid cell at three magnitudes spaced 1 mag apart, with the center magnitude at the approximate magnitude of the JAGB determined after iteration (for instance, for NGC 1448 stars are placed at 23.35, 24.35, and 25.35 mag in F150W); the central magnitudes for F150W and F277W are listed in Table 5. This results in a total of 600 stars per grid cell. We then calculate the mean difference between the input and output magnitudes and fit this bias as a function of input magnitude using an unweighted linear least squares fit. For each star, we use this fit to solve for the ‘true’ magnitude before bias, that is,
where is the magnitude of the JAGB stars after crowding correction, is the observed magnitude of the same star, is the slope of the fit between the crowding bias and magnitude, and is the intercept of that same fit. We use the Rockfish cluster at Advanced Research Computing at Hopkins (ARCH) together with GNU parallel (Tange, 2011) to significantly reduce the time needed to run these artificial star measurements. In addition to determining the crowding bias, we also use the artificial star test results for NGC 5584, the furthest of the galaxies analyzed here, to test for possible incompleteness of the photometry. At the faintest test point of 1 mag fainter than 24.9 mag (25.9 mag) in F150W for NGC 5584 we recover approximately 97% of artificial stars in the region used to measure the JAGB. We provide the photometry used to measure the JAGB after crowding correction and crossmatching in the Github repository (upon publication), https://github.com/siyangliastro/JWST_JAGB_GO1685, and the CMDs using this photometry in Fig. 5.
| Galaxy | Visit | Module | Input F150W | Input F277W | Mean Crowd F150W | Mean Crowd F277W |
| NGC 4258 | 1, 2 | B, A | 22.5 | 21.3 | 0.04 | 0.04 |
| NGC 1448 | 1, 2 | B, A | 24.35 | 23.05 | 0.02 | 0.01 |
| NGC 1559 | 1, 2 | B | 24.4 | 23.2 | 0.09 | 0.06 |
| NGC 5584 | 1, 2 | B | 24.9 | 23.7 | 0.06 | 0.04 |
| NGC 5643 | 1, 2 | A, B | 23.55 | 22.35 | 0.04 | 0.04 |
2.4 Selection of a JAGB Reference Magnitude
Perhaps the most challenging issue, based on the study of past efforts, is to define the JAGB standard candle, i.e., the procedure by which we assign a reference magnitude and its uncertainty to represent the apparent clumping of stars in the JAGB regions of the NIR color-magnitude diagram. We start with the example of the CMDs in NGC 4258, an important reference galaxy for distance measurements, to illustrate possible procedures.
In Fig. 4, we show CMDs for NGC 4258 containing the J-region in F150W vs. (see Fig. 3) on the West (chips A3+A4) and East (chips B3+B4) sides of the galaxy. We will refer to these two regions as the ‘Inner’ regions, as they appear to be on the inner disk and they are inside the “Outer” chips, A1, A2, B1, B2, at or near the halo. JAGB studies in the literature typically apply color and magnitude cuts to isolate the enhanced density of stars in the J-region (see, for instance, Zgirski et al. 2021); we approximate past efforts to put a “box” around the JAGB clump by defining a color cut of of 1 to 1.5 mag, and for NGC 4258 we use magnitude cuts of F150W = 21 to 24 mag (spanning about 5 standard deviations, i.e., a broad frame). This is the region marked as a blue dashed box in Fig. 4. We enlarge the points in the blue dashed box for emphasis. A binned luminosity function corresponding to stars inside the box can be seen in the right hand side of each subplot.
In the literature, the reference magnitude has been measured from this luminosity function in a variety of ways, such as from its mode, median, mean, or from a model, in some cases with additional, unstated specifications which thus make it difficult for us employ here. For example Madore & Freedman (2020) and Freedman & Madore (2020) use a mean but do not specify its range in magnitude or color a priori; rather the region of the CMD to average is chosen on a case-by-case basis from inspection of each CMD (based on the variety of selection boxes presented in their CMDs). As a result, the width in magnitude varies from 0.8 to 2.0 mag, and the red color boundary from 1.8 to 2.2.
To illustrate a specific prescription, we can use the mode-based method from Lee et al. (2023), which bins the luminosity function in 0.01 mag widths, applies Gaussian-windowed, locally weighted scatterplot smoothing (GLOESS; Loader, 2004; Persson et al., 2004) to the binned luminosity function with a smoothing scale of 0.25 mag, and uses the mode of the smoothed luminosity function to define the reference magnitude of the JAGB with an uncertainty set to the standard deviation of the LF divided by the square root of the number of stars in contains. We show the smoothed luminosity function for these measurements in blue and the corresponding mode-based JAGB measurement with the red line and label in Fig. 4.
Unfortunately, the luminosity functions we observed for the JAGB in NGC 4258 are quite asymmetric (skew 0.4) which makes the selection of the measure of the reference magnitude non-trivial because its value will differ substantially depending on the choice of method, and this choice will not cancel when we compare to other LFs if their shapes are not the same. The asymmetry seen here is not a fluke or the result of incompleteness or extinction; in fact, a varying degree of one-sided asymmetry appears to be intrinsic to JAGB luminosity functions. As previously noted, asymmetric examples for nearby galaxies include the LMC and NGC 6822 (skew 0.3-0.5), while symmetric examples include the SMC and IC 1613 (skew , Parada et al., 2021).
Depending on the level of LF asymmetry, we note that even the measurement of its mode will vary depending on the smoothing scale used as illustrated in the Appendix A, giving us an additional choice to consider. It is also far from obvious which measure of the LF will yield the better standard candle; this depends on whether the breadth of the LF contains multiple populations, the most common of which may be the standard candle, or whether it results from the spreading of a population whose mean is the true candle. Literature studies generally choose one method but none have provided a systematic study demonstrating the superiority of one or another.555Although as previously noted, the mean and median are preferred over the mode for comparing JAGB and the geometric distances for the LMC and SMC which we review in the Discussion. Therefore we will explore several different measures and retain a “many methods” approach to comparing these.
We use the following four measures of the JAGB LF: median, 3-sigma clipped mean, mode, and GaussianQuadratic model fit as used (or modified) from recent JAGB literature (Madore & Freedman, 2020; Freedman & Madore, 2020; Ripoche et al., 2020; Zgirski et al., 2021; Parada et al., 2021; Lee et al., 2021b, 2022, a; Parada et al., 2023; Lee, 2023). We use 3 clipping to the mean to reduce sensitivity to outliers and box boundaries, features already inherent with the median and mode. In addition, we use the simple median instead of the modified median combined with skew from Ripoche et al. (2020); Parada et al. (2021, 2023). We will also vary the color range used to select the J-region stars in five different ways. In all we will use 25 combinations (5 measures, 5 color ranges) as listed in Table 6 to provide a comprehensive view.
Ideally we seek a standard candle which is least sensitive to variations in underlying, intrinsic properties we cannot (yet) easily measure, such as age and metallicity; therefore we proceed to assess how the 25 possible combinations of JAGB measures vary in different fields of the same host.
3 JAGB Reconnaissance: Tests within Hosts
We first define a reference JAGB LF in the outer disk of each host before proceeding to divide the measures within a host for testing. For each host, we select stars with the same color cuts described in Section 2 (1 to 1.5 mag) and outside ellipses which contain the main disks of the host (see Figure 3) with parameters and magnitude cuts listed in Table 7. We discuss these choices of ellipse parameters and the effect of the measured JAGB with radius in Appendix B. We plot the CMDs for all host galaxies in Fig. 5 together with the luminosity function in the J-region and its mode-based magnitude; we also list the differences between mode and mean, the skewness, and the number of stars in that region.
We can test the stability of the JAGB by comparing fields around each host that are relatively equidistant from their main disks. The observation orientations for each host (see Figure 3) provide for such a test by comparing the fields 180 degrees apart for NGC 4258, NGC 5643 and NGC 1448. For NGC 5584 and NGC 1559 we will compare fields around the circumference of each disk.
| Variant | Range Considered | ||||
| Method | Median | 3 clipped mean | Mode | Gaussian+Quadratic Model Fit | |
| Smoothing | 0.25, 0.35, 0.4 | ||||
| Color Range | Narrow | Baseline | Wide | Blue 5% | Red 5% |
| Both sides 5% narrower | Both sides 5% wider | Blue side 5% redder | Red side 5% bluer |
3.1 NGC 4258
We first compare measured JAGBs in the East vs. West fields on opposite sides of NGC 4258 (see footprints in Fig. 3). We construct three cases for this comparison using different combinations of chips: Outer (A1, A2 vs. B1, B2), Inner (A3, A4 vs. B3, B4), and Inner+Outer (Full A module vs. Full B module).
The CMDs for the Inner case can be found in Fig. 4, and we plot the CMDs for the Outer and Inner+Outer cases in Fig. 6. From these figures, we can immediately see an asymmetry (skew 0.4, similar to the LMC) in the J-region luminosity functions at a level which produces a difference of 0.3 mag between the mode and mean, as shown in the upper right hand corners of each subplot. We also plot the differences between the measured JAGBs in the East and West fields for all three cases, together with measurement and selection variants, in Fig. 7. For the Outer fields model fits only, the sparsity of stars makes it difficult to achieve convergence and so we bin the luminosity function with widths of 0.2 mag instead of widths of 0.1 mag. From Fig. 7, we find that the Outer fields show variations between the sides of NGC 4258 on the order of 0.1-0.2 mag for all methods; these are larger than their statistical uncertainties, indicating that the differences are intrinsic. The Inner fields show better agreement, although agreement to within their statistical uncertainties depends on measurement and selection choices. It was noted by Lee et al. (2023) that the JAGB magnitude may converge at greater radii from the host, but this may run counter to the sense here of the better agreement for the Inner fields. In the interest of greater field-to-field averaging, we will use the combined Inner+Outer fields from both sides to calibrate other host JAGB measurements (see Section 4 for more discussion).
3.2 NGC 1448 & 5643
Next, we repeat the symmetry null test described in Section 3.1 for NGC 1448 and NGC 5643. For NGC 1448, we combine photometry from visit 1, module B and Visit 2, module A, and average the magnitudes for overlapping stars. Following the recommendation of using only stars in the outer disk for JAGB measurements from Lee et al. (2022) and Lee (2023), we select stars in the outer disk that lie outside the “disk” ellipse with parameters listed in Table 7; we then further split the outer disk region into the Northwest and Southeast regions, as labeled in Fig. 3. For NGC 5643, we compare the North and South modules (visit 2 module B vs. visit 1 module A) as labelled in Fig. 3. We plot the CMDs in Fig. 8; the differences in measured JAGBs between the Northwest and Southeast fields for NGC 1448 and the North and South fields for NGC 5643 are shown in Fig. 8. We find that the region differences in the apparent JAGB reference magnitudes between the two fields within NGC 1448 and NGC 4258 are centered at approximately mag and mag, respectively, while the JAGB reference magnitudes for NGC 5643 mostly agree to within one sigma and generally 0.05 mag, suggesting that the consistency of the JAGB is better in some galaxies compared to others.
3.3 Angular Symmetry Tests
Because the measured JAGBs in Section 3 did not all pass the test of consistency between 180 degree opposite fields, we decided to further investigate their variation with angle around NGC 1559 and NGC 5584, whose outer disks are well contained within the NIRCam frame. We do not use NGC 1448 or NGC 5643 because they are larger in size such that the GO-1685 observations do not include the perimeters of their outer disks.
We first combine the two epochs of observations for NGC 1559 and NGC 5584 and average stars that are observed twice. We then define a window 90 wide beginning with a position angle of 0, where 0 corresponds to the unit coordinate of (1, 0). We then slide each window by 10 and repeat the JAGB measurement until lower bound of the window edges reaches 350. We repeat this test for each galaxy using four different representative measurement methods (mode with 0.25 and 0.35 mag smoothings, median, and sigma-clipped mean). We do not show the model fit as the mode with mag behaves very similarly. In addition, we choose these methods to best approximate the methodological choices from the literature (for instance, from Madore & Freedman, 2020; Ripoche et al., 2020; Lee et al., 2023). We also impose a minimum threshold of 100 stars to minimize fluctuations due to low number statistics and plot the results in Fig. 10. Yellow stars mark windows that do not overlap and are uncorrelated. Statistical uncertainties are calculated using the error on the mean. We observe differences of 0.1 mag that are not consistent with statistical uncertainties and persist across measurement methods, again suggesting some intrinsic differences in the measured JAGB.
From the tests described above, we find evidence of spatial variations in the JAGB reference magnitude. These variations depend on the method used to fix the JAGB reference magnitude and may be reduced by observing a larger portion of the outer disks. We discuss these variations in a broader context in Section 5.
| Galaxy | Semi-Major Axis (SMA) | Minor/Major Axis Ratio | PA | Cuts |
| NGC 1448 | 200 | 0.224 | 41 | 23 to 25.5 mag |
| NGC 1559 | 75 | 0.523 | 65 | 23 to 25.8 mag |
| NGC 5584 | 75 | 0.64 | 158 | 23.5 to 26 mag |
| NGC 5643 | 75 | 0.871 | 89 | 22.5 to 24.8 mag |
4 Distance Ladder and
In this section, we measure distances to NGC 1448, NGC 1559, NGC 5584, and NGC 5643 using the JAGB methods defined in the previous sections. For each measurement method, we compare the SN host JAGB magnitudes to those from the combined NGC 4258 Inner+Outer fields, which best approximates the outer disk, and use the geometric maser distance of = 29.397 0.032 mag from Reid et al. (2019) to calibrate the SN host distances. The JAGB magnitudes for NGC 4258 across all variants are listed in Table 8 and plotted in the leftmost panel of Fig. 11. The other four panels of Fig. 11 give the corresponding distance modulus to each of the four galaxies for all variants, with the average shown as a red horizontal line.
For reference, we also give the distance moduli obtained from Cepheids (blue), TRGB (yellow), and Miras (cyan) where available; the Cepheid distances are obtained from the solution in (Riess et al., 2022). To eliminate calibration differences we adjust these distances upwards by 0.02 mag to reflect the distance scale based on NGC 4258 as the sole anchor. The resulting Cepheid distance moduli are = 31.31 mag, = 31.51 mag, = 31.79 mag, and = 30.56 mag, with the solid blue lines. The Mira distance to NGC 1559 is mag (Huang et al., 2020). The TRGB distances of = 31.38 mag and = 30.47 mag are from the Extragalactic Distance Database (EDD) (Tully et al., 2009; Anand et al., 2021), while = 31.49 mag and = 31.80 mag are from JWST measurements from Anand et al. (ApJ, submitted).
We find broad agreement between JAGB distances and those based on the Cepheid, TRGB, and Miras, with the variation in the distances spanning approximately 0.2 mag. As expected, the mode is closer to the mean and median measures for the case of larger LF smoothing, likely due to the GLOESS smoothing removing asymmetry in the luminosity function. The model fit also tends to agree better with a higher smoothed mode, likely due to the Gaussian component of the model fit.
| Method | Color Selection | Smoothing | ||
| 22.539 | 0.027 | Median | Narrow | nan |
| 22.594 | 0.027 | 3 Clipped Mean | Narrow | nan |
| 22.372 | 0.027 | Mode | Narrow | 0.25 |
| 22.432 | 0.027 | Mode | Narrow | 0.35 |
| 22.419 | 0.044 | Model Fit | Narrow | nan |
| 22.568 | 0.024 | Median | Baseline | nan |
| 22.611 | 0.024 | Clipped Mean | Baseline | nan |
| 22.392 | 0.024 | Mode | Baseline | 0.25 |
| 22.452 | 0.024 | Mode | Baseline | 0.35 |
| 22.445 | 0.042 | Model Fit | Baseline | nan |
| 22.606 | 0.023 | Median | Wide | nan |
| 22.641 | 0.023 | Clipped Mean | Wide | nan |
| 22.441 | 0.023 | Mode | Wide | 0.25 |
| 22.511 | 0.023 | Mode | Wide | 0.35 |
| 22.516 | 0.034 | Model Fit | Wide | nan |
| 22.539 | 0.027 | Median | Blue 5 | nan |
| 22.595 | 0.027 | Clipped Mean | Blue 5 | nan |
| 22.372 | 0.027 | Mode | Blue 5 | 0.25 |
| 22.432 | 0.027 | Mode | Blue 5 | 0.35 |
| 22.424 | 0.042 | Model Fit | Blue 5 | nan |
| 22.568 | 0.025 | Median | Red 5 | nan |
| 22.61 | 0.025 | Clipped Mean | Red 5 | nan |
| 22.382 | 0.025 | Mode | Red 5 | 0.25 |
| 22.452 | 0.025 | Mode | Red 5 | 0.35 |
| 22.442 | 0.044 | Model Fit | Red 5 | nan |
Next, we average over the four hosts and compare the median JAGB distances and Cepheid-based distances as shown in Fig. 12 for each measurement variant. This sample difference allows us to determine the value of relative to the complete distance ladder provided in Riess et al. (2022), i.e., allowing for this single, measured offset between this sample of differences between the first and second rungs, .
Unfortunately, as seen in Fig. 12, this sample difference does depend quite significantly on which measurement method we choose, with a full range of 0.18 mag. In the next section we identify the source of this methodological uncertainty as originating from the varying degree of asymmetry of the JAGB LFs. All of the luminosity functions exhibit to varying degrees a positive skew and negative difference between the mode and mean-based JAGB measurements, as shown in Fig. 5. The fact that the calibrator, NGC 4258, has the largest skew (but not unusual in degree as it is comparable to that of the LMC and NGC 6822), with a mean minus mode of 0.29 mag (versus the mean SN host with mean minus mode of 0.1 mag), produces a systematic bias that depends on the measure employed.
Despite the method-dependent bias, the JAGB observations in the anchor NGC 4258 and four SN Ia host galaxies still provide interesting constraints on the Hubble constant. We may adopt a fiducial result and uncertainty from the median of the differences between the JAGB and Cepheid distances as shown in Fig. 12 to derive relative to that in Riess et al. (2022), i.e. 5log( / ) = (Cepheids - JAGB). Calculating relative to that from Riess et al. (2022) allows us to establish a measurement using the same treatment of SN Ia from Riess et al. (2022) and isolate differences in due to only differences in JAGB and Cepheid distances between rung one and two. We find (stat) 2.3 (sys) ( 3.2 combined) km s Mpc with a range from 72 78 km s Mpc from variations in the choice of measurement method. We list the error budget in Table 9. In the Discussion we consider whether there is compelling evidence to select any specific method over the spread of methods as to indicate a better result.
| Error | Value (stat) [mag] | Value (sys) [mag] |
| NGC 4258 Maser Distance | 0.03 | |
| Methodological Variations | 0.06 | |
| JAGB Mean four host Measurement | 0.04 | |
| Cepheid Mean four host Measurement | 0.04 | |
| Rest of DL: SN Ia, local and Hubble Flow | 0.025 | |
| Total | 0.06 | 0.07 |
5 Discussion
We investigated the JAGB using observations of NGC 4258, NGC 1448, NGC 1559, NGC 5584, and NGC 5643 from GO-1685 (P.I.: A. Riess, Riess et al., 2021), testing the degree of internal consistency of the measurements from different subsets of the JAGB samples. A series of symmetry and angular uniformity tests yield clear evidence of variations in JAGB luminosities larger than their measurement uncertainties and confirms past evidence (Parada et al., 2021) of differences in the degree of asymmetry of their luminosity functions. We further verified that the LF asymmetry is that of the host stars in the JAGB selection box and not due to contamination from background galaxies which are well-resolved by JWST and identified by DOLPHOT’s star/galaxy discriminator (with compact, background galaxies also generally too blue to appear in the selection box). We also note that foreground stars are also far too sparse to contaminate the host CMDs, as a similar field size of the Hubble Deep (or Ultra Deep) Field at similar galactic latitude as these hosts yields only O() stars in the relevant magnitude range of , and fewer in the selected JAGB color range.
Our findings are supported by the results of Parada et al. (2021), who had also found both variations in the JAGB luminosity and differences in the shape of J-region luminosity functions via skew, with the LMC and NGC 6822 being more asymmetric than the SMC or IC 1613. Likewise, Lee et al. (2022) identified variations in the asymmetry and luminosity of the JAGB LF in M33 which varied with radial distance from the host with the LF being most symmetric in a “sweet spot” (region 3, between regions 1, 2 and 4). In Appendix B, we present further study of the radial differentiation of the JAGB.
In Fig. 13 we compare the skewness of the JAGB luminosity functions examined in this paper by fitting them with the Lorentzian model described in Parada et al. (2021). We also include the LF for the LMC and SMC with H-band photometry retrieved from Macri et al. (2015) and Ripoche et al. (2020), respectively. The legend provides the Fisher-Pearson skew and the difference between the JAGB measured using the mode of the fit and the 3 clipped mean. We offset the LFs in magnitude based on their modes (left panel) or their sigma-clipped mean (right panel); the difference clearly illustrates how mode- and mean-based measurements differ for skewed LFs. We note that the band JAGB can also be skewed by similar amounts, as shown in Lee (2023); the same holds for the J-band LF of the LMC (Parada et al., 2021). This shows that the skew effect is not unique to a particular NIR band. In this comparison, we use the Lorentzian fit to the LFs, because GLOESS smoothing tends to naturally reduce the LF skew; this can be seen in the Appendix A. As a result, a larger smoothing window used before calculating the mode produces a result closer to the mean and median measurements (see for instance, Fig. 12). The source of variation in inferred primarily comes from the difference in the skews between NGC 4258 and host galaxies. In NGC 4258, the difference between the mode and mean is 0.29 mag, while in the host galaxies, it is 0.1 mag resulting in a 0.2 mag difference across the distance ladder.
5.1 Asymmetry in the Clouds
It is instructive to further scrutinize the JAGB LFs in the Magellanic Clouds for further guidance, as these exhibit a disparity in asymmetry similar to our target galaxies, while having the virtue that their relative distances are well-established geometrically and they have been well studied in the literature using various measurement methods. In Table 10 we provide JAGB reference magnitudes in the LMC and SMC taken from the literature, together with our measurements using the same suite of methods used in our analysis (i.e., mode with , sigma-clipped mean and the median) on the same datasets as prior authors. For the LMC, which has an asymmetric JAGB LF (Parada et al. (2021) reports a skew of ), past studies of the J vs CMD from several authors (Madore & Freedman, 2020; Freedman & Madore, 2020; Ripoche et al., 2020; Parada et al., 2021; Zgirski et al., 2021) give an apparent reference magnitude of , a range of 0.15 mag; the difference between these results is much larger than the typical quoted statistical uncertainty of 0.004 mag, which is derived from the standard deviation of the distribution (0.2–0.3 mag) divided by the square root of the number of stars (typically several thousands). With our measurements, we similarly find that the asymmetry produces a large, method-dependent range of 0.1 mag. The choice of smoothing for the mode-based method also slightly shifts the JAGB reference magnitude. Another example of this effect can be seen in Appendix A. The SMC, which has a symmetric luminosity function, has a far more stable JAGB reference magnitude that also does not change with different smoothing values. We find that the mean and median for the J-region luminosity function using the Macri et al. (2015) differ by 0.06 mag.
We can check which method provides a more standardized reference magnitude by comparing the differences between the geometric distances and JAGB reference magnitudes between the SMC and LMC from Graczyk et al. (2020) and Pietrzyński et al. (2019), restrict the SMC sample to stars within 1 degree within the SMC center to cover the same region as the DEBs while avoiding the complexity of the depth of the SMC outside its core region. We measure the JAGB reference magnitudes using the same variants described above and plot the differences between the SMC and LMC reference magnitudes in Fig. 14. We shade in gray the area corresponding to the difference between the SMC and LMC geometric distances of 0.500 0.017 mag. We find that in general, the median and clipped mean methods agree most closely with the geometric differences; these methods also tend to produce the least variation in the null and angular tests shown in Section 3.3. Methods which try to measure the peak of the LF, such as the mode (especially with little LF smoothing) or a model fit to the peak, are in significantly larger disagreement with the geometric difference.
Non-uniformity in the color-dependence of the luminosity function can also introduce some variation in the standard candle. In all JWST host galaxies we found that the reference magnitude varies somewhat with color, not linearly, but tending to be brighter near the middle of the color range. If the color dependence is uniform across hosts, it will have no impact on distance measurements determined by comparing different hosts. In practice, the relation between the reference magnitude and color appears less important than the shape of the luminosity function based on our comparisons of different color ranges.
| Galaxy | Study | Sample | Method | Ref. | Anchor Dist | Anchor Ref. | |||
| (1) | (2) | (3) | (4) | (5) | (6) | (7) | (8) | (9) | (10) |
| LMC | MF20 | M15 | Selected range mean | 12.31 | 0.05 | NED | 18.477 | P19 | -6.22 |
| LMC | Here | M15 | Median | 12.37 | 0.08 | S21 | 18.477 | P19 | -6.21 |
| LMC | Z21 | K09 | Model Fit | 12.38 | 0.11 | G20 | 18.477 | P19 | -6.21 |
| LMC | R20 | R20 | ML Median | 12.31 | 0.11 | G20 | 18.477 | P19 | -6.28 |
| LMC | P21 | R20 | Mode (Lorentzian model) | 12.23 | 0.11 | G20 | 18.477 | P19 | -6.36 |
| LMC | Here | R20 | Median | 12.31 | 0.08 | S21 | 18.477 | P19 | -6.25 |
| LMC | Here | R20 | Mode = 0.25 | 12.29 | 0.08 | S21 | 18.477 | P19 | -6.27 |
| LMC | Here | R20 | Mode = 0.35 | 12.31 | 0.08 | S21 | 18.477 | P19 | -6.26 |
| SMC | MF20 | S06 | Selected range mean | 12.81 | 0.03 | NED | 18.965 | G20 | -6.18 |
| SMC | Z21 | K09 | Model Fit | 12.85 | 0.07 | G20 | 18.977 | S16 | -6.20 |
| SMC | R20 | R20 | ML Median | 12.87 | 0.07 | G20 | 18.96 | S16 | -6.16 |
| SMC | P21 | R20 | Mode (Lorentzian model) | 12.90 | 0.07 | G20 | 18.96 | S16 | -6.15 |
| SMC | Here | R20 | Median | 12.83 | 0.04 | S21 | 18.977 | G20 | -6.19 |
| SMC | Here | R20 | Mode = 0.25 | 12.85 | 0.04 | S21 | 18.977 | G20 | -6.17 |
| SMC | Here | R20 | Mode = 0.35 | 12.85 | 0.04 | S21 | 18.977 | G20 | -6.17 |
5.2 The JAGB in Perspective
We hypothesize that these variations in the shapes of the JAGB luminosity function may result from intrinsic, astrophysical differences in the distribution of carbon star demographics. Parada et al. (2021) suggested that some of the field-to-field variations in the measured JAGB luminosity seen in the LMC, listed in their Table 5 and described in their Fig. 5.1, may be due to variations in the star formation history in different regions. The differences are unlikely to result from interstellar reddening, due to the use of reddening maps to remove absorption. We would also not expect interstellar reddening to play a role in the host studied here, based on estimates of the variation in interstellar reddening with radial separation presented by Ménard et al. (2010) (see equation 29). For example, at semi-minor axes separations of 39 and 48 from the centers of NGC 1559 and NGC 5584 corresponding to the ellipses in Table 7, we expect NIR internal extinctions of 6 mmag and 5 mmag, respectively. Parada et al. (2021) also suggest that metallicity can drive the skewness of the J-region luminosity function and the JAGB luminosity, and choose either to use the LMC or SMC as anchors based on the skewness of the J-region luminosity functions in their target galaxies. We also note that it may also be possible that the J-region contains a mixed population of stars, such that not all groups of stars have the same mean magnitude. Further work is needed to investigate the origin of the variations in the measured JAGB and to demonstrate that these differences in treatment of different targets do not create undue variations and uncertainties in the estimated relative distances. Because we do not know whether the often asymmetric shape of the LF is due to the spread of a single population or the combination of multiple symmetric populations, it is not clear a priori which reference flux should provide the best standard candle, something that is crucial to take into consideration when using the JAGB as a standard candle.
The investigations of variations in the JAGB presented above highlight several issues to consider when using the JAGB to measure distances. It is evident that the shape of the JAGB luminosity function, as it is currently selected, is non-uniform and often asymmetric . This produces systematic uncertainties in distance measurements from the data presented here (and in the Clouds) resulting from the different and credible options of defining a reference flux. We do not think it is plausible to select just one method or define a somewhat different measurement prescription or magnitude range for each host and ignore the uncertainty arising from the freedom in these choices.
For an overall perspective, we plot in Fig. 15 the JAGB distances measured here against the Cepheid distances from Riess et al. (2022), adjusted for a NGC 4258 only anchor, together with the JAGB and Cepheid distances from Zgirski et al. (2021). For our distance measurements we use the weighted average of the distances shown in Fig. 11, and include the contribution of methodological variation in the estimated errors. We find that the JAGB and Cepheid distances show reasonable agreement within their error bars.
The JWST data presented above demonstrate that until JAGB LF variations are standardized or understood, they should be included in the uncertainties when the JAGB is used to measure extragalactic distances and the Hubble constant. At present, this issue lay in the critical path to reach the few percent level precision in measuring the Hubble constant.
Such issues are not unexpected for the development and refinement of a new standard candle. Recent work has already made inroads into improving the choice of radius at which to measure JAGB (Lee, 2023) (see discussion comparing radii in Appendix B) as well as how best to match the shape of the luminosity function to measure distances (Parada et al., 2023). Future studies and JAGB observations are likely to continue improving this promising standard candle.
Appendix A Effects of Smoothing on J-region Luminosity Functions
We find that the choice of smoothing in the mode-based method described in Lee et al. (2022); Lee (2023) can bias JAGB measurements for asymmetric luminosity functions. We demonstrate this by plotting in Fig. 16 the smoothed luminosity function for NGC 4258, Inner+Outer West field in F150W using baseline color cuts of 1 mag and 1.5 mag, magnitude cuts of 21 mag and 24 mag, and three levels of smoothing ( 0.2, 0.3, and 0.4 mag). We note that smoothings of 0.2 mag is used by Lee (2023) and 0.25 and 0.3 mag are used by Lee et al. (2023). The difference between the brightest and faintest measured JAGBs for the range of smoothing from 0.2 and 0.4 mag is 0.09 mag. We caution that using different smoothing parameters with this measurement method across the distance ladder can potentially introduce biases if luminosity functions are not equally asymmetric in the anchor and host galaxies.
Appendix B Stability of the JAGB with Radius
Lee et al. (2022) and Lee et al. (2023) show that the measured JAGB can change with radius in a galaxy and argue that the JAGB should be measured in a region where the measured JAGB converges. To find where the measured JAGB converges, Lee et al. (2023) construct annuli centered on their target galaxy and plot the measured JAGB as a function of the DOLPHOT sky parameter in each annulus.
We perform a similar test with our sample by constructing several annuli centered on the galaxy and measure the JAGB using the stars inside each annulus. We use the same color and magnitude cuts as described in the main body of this paper. For the purpose of this exercise, we used the mode based method and smoothing from Lee et al. (2022); Lee (2023) to measure the JAGB.
For NGC 1448, we begin constructing annuli with semimajor axes (SMA) of 80SMA100 and increase each of these boundaries by 20 until we reach 200SMA220. For NGC 1559 and 5584 we begin with 20SMA40 and increase each boundary in 20 increments until we reach 140SMA160. We plot the mean DOLPHOT sky value as a function of the midpoint annulus SMA and the corresponding measured JAGBs as a function of the midpoint annulus SMA in Fig. 17. We mark with the red line the SMA used to construct the outer disk regions in the main body of this paper for reference. We find that not all galaxies show measured JAGBs that converge out to a large radius (for instance, the measured JAGBs for NGC 5584 appear to become increasingly fainter to the left of the red line without indication flattening in the same way with NGC 1559) suggesting that there may need to be additional standardization beyond a radial cut. We note that convergence does not necessarily imply that the JAGB populations across galaxies are similar and yield the same intrinsic JAGB luminosity. For instance, convergence may occur at regions of differing metaliicities which could change the JAGB luminosity (see Section 1 for a discussion on a possible metallicity trend).